OpenStudy (x3_drummerchick):

Help! will give medals if you can explain: A soccer ball m=0.65kg is dropped from height h=4.9m bounces off the grass. The ball loses E=3.4J of energy to the environment during the bounce. Write an expression for high (final height) the soccer ball bounces.

OpenStudy (x3_drummerchick):

@TheSmartOne @IrishBoy123

OpenStudy (x3_drummerchick):

@.Sam. @Elsa213

OpenStudy (irishboy123):

Initial Energy (Potential) : \(U_1 = mgh_1\) When it hits the grass, Energy (Kinetic) is : \(T_1 = \frac{1}{2} m v_1^2 = U_1\) Energy immediately after collision (Kinetic): \(T_2 = T_1 - 3.4 = U_1 - 3.4\) Once it rises and converts all that kinetic into potential energy, we say that: \(U_2 = T_2 = U_1 - 3.4\) \(\implies mgh_1 = mgh_o - 3.4\) \(\implies h_1 = h_o - \dfrac{3.4}{mg}\) \(h_1 = h_o - \dfrac{3.4}{0.65 * 9.81}\) m

OpenStudy (x3_drummerchick):

@IrishBoy123 , what is the initial expression you used for h final? h Final= h initial-(E/(mg))?

OpenStudy (x3_drummerchick):

\[ho-(\frac{ E }{ mg })m\] keeps coming up as incorrect?

OpenStudy (irishboy123):

Yes \(h_f = h_i - \dfrac{\mathbf{Energy loss}}{0.65 * 9.81}\) m Initial Energy (which is Potential) : \(E_i = U_i = mgh_i\) When it hits the grass, Energy (which is Kinetic) is : \(E_i = T_i = \frac{1}{2} m v_i^2 = U_i\) Energy immediately after collision (Kinetic): \(E_f = T_f = T_i - 3.4 = U_i - 3.4\) Once it rises and converts all that kinetic into potential energy, we say that: \(E_f = U_f = T_f = U_i - 3.4\) \(\implies mgh_f = mgh_i - 3.4\) \(\implies h_f = h_i - \dfrac{3.4}{mg}\) \(h_f = h_i - \dfrac{\mathbf{Energy loss}}{0.65 * 9.81}\) m

OpenStudy (irishboy123):

oh, :-( the m is metres, so sorry

OpenStudy (irishboy123):

\(h_f = h_i - \dfrac{\mathbf{Energy loss}}{0.65 * 9.81}\) putting the unit in.... \(h_f = \left( h_i - \dfrac{\mathbf{Energy loss}}{0.65 * 9.81}\right) \mathbf{m}\) mea culpa :(

OpenStudy (x3_drummerchick):

i messed up with my paeantheses, thank you!!! i had it right the whole time, i just never used the correct placement of parantheses, thnak you!

OpenStudy (x3_drummerchick):

if i could give you a 2nd medal, I would. thank you :)

OpenStudy (irishboy123):

lol!! :)